.A Corporation sold 5000 issue of 20-year bonds, having a total face value of 10,000,000 for 9,500,000. The bonds bear interest at 16%, payable semiannually. The company wishes to establish a sinking fund for retiring the bond issue and will make semiannual deposit that will earn 12%, compounded semiannually. Compute the annual cost for interest and redemption of these bonds. If an investor wishes to dispose of his 100 shares of his bonds at the end of the tenth year, at what price must he sell his bond to give him a profit of 15% nominal on his investment?
Answer #1

Nominal value per bond = 10000000/5000

= 2000

Person holding 100 shares of bonds nominal value. = 2000× 100 = 200000

Profit of 15% = 30000

Purchase value of 100 bonds = 9500000/5000*100 = 190000

To get this profit he has to sell at value of

220000